jigglypuff314 (jigglypuff314):

What will be the remainder when 3^99 is divided by 5?

OpenStudy (anonymous):

3^2 mod 5 = 4 3^3 mod 5 = 2 3^4 mod 5 = 1 3^5 mod 5 = 3 3^6 mod 5 = 4 3^7 mod 5 = 2 3^8 mod 5 = 1 3^9 mod 5 = 3

OpenStudy (jdoe0001):

hmmm I ended up with a 7

OpenStudy (jdoe0001):

or so the calculation said =)

jigglypuff314 (jigglypuff314):

okay, thank you all! :)

OpenStudy (jdoe0001):

yw

OpenStudy (raden):

3 ^ 99 = 3^3 * 3^96 = 3^3 * (3^4)^24 = 3^3 * 81^24 = ....7 if a big number which has the unit digit is 7 when divided by 5 then its remainder is 2